OpenStudy (anonymous):

what is the ? in: (8x/x+5)(?/24x)=1/3

OpenStudy (anonymous):

please help me out

mathslover (mathslover):

is it \[\huge{\frac{x}{x+5}}\]

OpenStudy (anonymous):

@mathslover how did you get the answer??

mathslover (mathslover):

i didn't get the answer i am asking that is it 8(x/(x+5)) or 8((x/x)+5) ?

OpenStudy (maheshmeghwal9):

\[\frac{8x}{x+5} \times \frac{?}{24x} \implies \frac{8x}{24x} \times \frac{?}{x+5} \implies \frac{1}{3} \times \frac{?}{x+5}=\frac{1}{3}\]So \[? = x+5.\]

OpenStudy (maheshmeghwal9):

becoz \[\frac{\cancel {x+5}}{ \cancel {x+5}}=1.\]

OpenStudy (maheshmeghwal9):

Have u any doubt?

OpenStudy (anonymous):

i just have a question,--is 1 the value of ?

OpenStudy (maheshmeghwal9):

no (x+5) is the value of ? becoz if u put this value @ ? then see what do u get; & tell me.

OpenStudy (anonymous):

owwwww, sorry for the doubt

OpenStudy (maheshmeghwal9):

no worries:)

OpenStudy (anonymous):

sorry again